# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ B.3 $b_{3}$ term
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.31)
@@
$\mathcal{V}_{b_{3}}=-\frac{k_{1}^{2}-k_{2}^{2}+k_{3}^{2}}{2k_{1}^{2}}\mathcal{V}_{b_{1}},$ (B.32)
@@
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (B.33)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (B.34)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (B.35)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.36)
In the case of (i), we obtain
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}),$ (B.37)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(-k_{1},k_{2},k_{3}),$ (B.38)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (B.39)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (B.40)
@@
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}}.$ (B.41)
In the other three cases, $\mathcal{I}_{b_{3},K^{\prime}}$ and $\mathcal{I}_{b_{3},K_{1}^{\prime}}$ take the same form with Eq. (B.37) and Eq. (B.38), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (B.42)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (B.43)
@@
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}).$ (B.44)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
@@ B.4 $b_{4}$ term
$\quad\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.45)
@@
$\mathcal{I}_{b_{4},K^{\prime}}$ $:=-i\int{\rm d}\eta\eta^{2}e^{iK^{\prime}\eta},$ (B.47)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{1}^{\prime}\eta},$ (B.48)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{2}^{\prime}\eta},$ (B.49)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{3}^{\prime}\eta}.$ (B.50)
In the case of (i), we have
$\mathcal{I}_{b_{4},K^{\prime}}$ $=\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}),$ (B.51)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(-k_{1},k_{2},k_{3}),$ (B.52)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (B.53)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (B.54)
@@
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}):=\frac{2}{K^{\prime 3}}.$ (B.55)
In the other three cases, $\mathcal{I}_{b_{4},K^{\prime}}$ and $\mathcal{I}_{b_{4},K_{1}^{\prime}}$ take the same form with Eq. (B.51) and Eq. (B.52), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (B.56)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (B.57)
@@
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3}):=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2})$ (B.58)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
@@ B.5 $b_{5}$ term
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.59)
@@
$\mathcal{I}_{b_{5},K^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (B.61)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(-i+c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (B.62)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (B.63)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.64)
In the case of (i), we have
$\mathcal{I}_{b_{5},K^{\prime}}$ $=\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}),$ (B.65)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(-k_{1},k_{2},k_{3}),$ (B.66)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (B.67)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (B.68)
@@
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}):=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}.$ (B.69)
In the other three cases, $\mathcal{I}_{b_{5},K^{\prime}}$ and $\mathcal{I}_{b_{5},K_{1}^{\prime}}$ take the same form with Eq. (B.65) and Eq. (B.66), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (B.70)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (B.71)
@@
$\mathcal{S}_{b_{5}}(k_{1},k_{2},k_{3}):=e^{iK^{\prime}\eta_{0}}\biggl{\{}\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr{\}}.$ (B.72)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
@@ B.8 Field redefinition
$\dot{h}^{(s)}_{{\bf k}}\simeq ic_{h}^{2}k^{2}\eta^{2}Hh^{(s)}_{\bf k}.$ (B.77)
The quadratic-order terms in the field redefinition involve at least one time-derivative of the tensor perturbations, which means that those do not contribute to the bispectrum evaluated at the end of inflation $\eta\to 0$ (i.e., $\dot{h}^{(s)}_{{\bf k}}\to 0$). Therefore, the field redefinition always yields subleading contributions.
@@ Appendix C $\mathcal{V}_{b_{i}}$ at the squeezed limit
$\quad\ -\frac{1}{8}\left(s_{2}+s_{3}\right)^{2}\left[1-(\hat{n}\cdot\hat{q})^{2}\left(s_{2}^{2}-s_{2}s_{3}+s_{3}^{2}\right)\right]\frac{q^{2}}{k^{2}}+O((q/k)^{3}),$ (C.1)
@@
×$\quad\times\biggl{\{}\left(s_{2}+s_{3}\right)^{2}-2(\hat{n}\cdot\hat{q})(s_{2}-s_{3})(s_{2}+s_{3})\frac{q}{k}-\biggl{[}8+(1-8(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+(1-8(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}$
$\quad\quad\ +2s_{2}(s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3})\frac{q^{2}}{4k^{2}}\biggr{]}+O((q/k)^{3})\biggr{\}},$ (C.2)
@@
×$\quad\times\left\{\left(s_{2}+s_{3}\right)^{2}+\Big{[}s_{2}-4(\hat{n}\cdot\hat{q})^{2}s_{2}+s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3}\Big{]}(s_{2}+s_{3})\frac{q}{2k(\hat{n}\cdot\hat{q})}\right.$ (C.3)
$\quad\quad\ \left.-\Big{[}2+(1-2(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+2(\hat{n}\cdot\hat{q})^{2}s_{2}s_{3}-(1+2(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}\Big{]}\frac{q^{2}}{k^{2}}+O((q/k)^{3})\right\},$ (C.4)
$\mathcal{V}_{b_{4}}$ $=\mathcal{V}_{b_{1}},$ (C.5)
$\mathcal{V}_{b_{5}}$ $=q^{2}\mathcal{V}_{b_{1}},$ (C.6)
$\mathcal{V}_{b_{6}}$ $=\frac{1}{8}[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}+s_{3})^{2}-(\hat{n}\cdot\hat{q})[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}^{2}-s_{3}^{2})\frac{q}{8k}$ (C.7)
@@
$\quad+O((q/k)^{3}),$ (C.8)
$\mathcal{V}_{b_{7}}$ $=q^{2}\mathcal{V}_{b_{6}}.$ (C.9)

### v3

@@ B.3 $b_{3}$ term
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr$], (193)
@@
$\mathcal{V}_{b_{3}}=-\frac{k_{1}^{2}-k_{2}^{2}+k_{3}^{2}}{2k_{1}^{2}}\mathcal{V}_{b_{1}},$ (194)
@@
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (195)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (196)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (197)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta}.$ (198)
In case (i), we obtain
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}),$ (199)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(-k_{1},k_{2},k_{3}),$ (200)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (201)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (202)
@@
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}}.$ (203)
In the other three cases, $\mathcal{I}_{b_{3},K^{\prime}}$ and $\mathcal{I}_{b_{3},K_{1}^{\prime}}$ take the same form as in Eq. (199) and Eq. (200), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (204)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (205)
@@
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}).$ (206)
In case (ii), $\eta_{0}$ is replaced with $1/(c_{s}k_{1})$.
@@ B.4 $b_{4}$ term
$\quad\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr$], (207)
@@
$\mathcal{I}_{b_{4},K^{\prime}}$ $:=-i\int{\rm d}\eta\eta^{2}e^{iK^{\prime}\eta},$ (209)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{1}^{\prime}\eta},$ (210)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{2}^{\prime}\eta},$ (211)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{3}^{\prime}\eta}.$ (212)
In case (i), we have
$\mathcal{I}_{b_{4},K^{\prime}}$ $=\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}),$ (213)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(-k_{1},k_{2},k_{3}),$ (214)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (215)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (216)
@@
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}):=\frac{2}{K^{\prime 3}}.$ (217)
In the other three cases, $\mathcal{I}_{b_{4},K^{\prime}}$ and $\mathcal{I}_{b_{4},K_{1}^{\prime}}$ take the same form as in Eq. (213) and Eq. (214), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (218)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (219)
@@
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3}):=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2})$ (220)
In case (ii), $\eta_{0}$ is replaced with $1/(c_{s}k_{1})$.
@@ B.5 $b_{5}$ term
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr$], (221)
@@
$\mathcal{I}_{b_{5},K^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (223)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(-i+c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (224)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (225)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (226)
In case (i), we have
$\mathcal{I}_{b_{5},K^{\prime}}$ $=\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}),$ (227)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(-k_{1},k_{2},k_{3}),$ (228)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (229)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (230)
@@
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}):=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}.$ (231)
In the other three cases, $\mathcal{I}_{b_{5},K^{\prime}}$ and $\mathcal{I}_{b_{5},K_{1}^{\prime}}$ take the same form as in Eq. (227) and Eq. (228), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (232)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (233)
@@
$\mathcal{S}_{b_{5}}(k_{1},k_{2},k_{3}):=e^{iK^{\prime}\eta_{0}}\biggl\{\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr\}.$ (234)
In case (ii), $\eta_{0}$ is replaced with $1/(c_{s}k_{1})$.
@@ B.8 Field redefinition
$\dot{h}^{(s)}_{{\bf k}}\simeq ic_{h}^{2}k^{2}\eta^{2}Hh^{(s)}_{\bf k}$. (239)
The quadratic-order terms in the field redefinition involve at least one time-derivative of the tensor perturbations, which means that those do not contribute to the bispectrum evaluated at the end of inflation $\eta\to 0$ (i.e. $\dot{h}^{(s)}_{{\bf k}}\to 0$). Therefore, the field redefinition always yields subleading contributions.
@@ Appendix C $\mathcal{V}_{b_{i}}$ at the squeezed limit
$\quad\ -\frac{1}{8}\left(s_{2}+s_{3}\right)^{2}\left[1-(\hat{n}\cdot\hat{q})^{2}\left(s_{2}^{2}-s_{2}s_{3}+s_{3}^{2}\right)\right]\frac{q^{2}}{k^{2}}+O((q/k)^{3}),$ (240)
@@
$\quad\times\biggl\{\left(s_{2}+s_{3}\right)^{2}-2(\hat{n}\cdot\hat{q})(s_{2}-s_{3})(s_{2}+s_{3})\frac{q}{k}-\biggl[8+(1-8(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+(1-8(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}$
$\quad\quad\ +2s_{2}(s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3})\frac{q^{2}}{4k^{2}}\biggr]+O((q/k)^{3})\biggr\},$ (241)
@@
$\quad\times\left\{\left(s_{2}+s_{3}\right)^{2}+\Big[s_{2}-4(\hat{n}\cdot\hat{q})^{2}s_{2}+s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3}\Big](s_{2}+s_{3})\frac{q}{2k(\hat{n}\cdot\hat{q})}\right$. (242)
$\quad\quad\ \left.-\Big[2+(1-2(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+2(\hat{n}\cdot\hat{q})^{2}s_{2}s_{3}-(1+2(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}\Big]\frac{q^{2}}{k^{2}}+O((q/k)^{3})\right\},$ (243)
$\mathcal{V}_{b_{4}}$ $=\mathcal{V}_{b_{1}},$ (244)
$\mathcal{V}_{b_{5}}$ $=q^{2}\mathcal{V}_{b_{1}},$ (245)
$\mathcal{V}_{b_{6}}$ $=\frac{1}{8}[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}+s_{3})^{2}-(\hat{n}\cdot\hat{q})[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}^{2}-s_{3}^{2})\frac{q}{8k}$ (246)
@@
$\quad+O((q/k)^{3}),$ (247)
$\mathcal{V}_{b_{7}}$ $=q^{2}\mathcal{V}_{b_{6}}.$ (248)